Calculate the root mean square velocity, in m/s, of carbon dioxide molecules at 100.0ºC.


1
Expert's answer
2021-01-20T03:16:47-0500

Molar mass of CO2,

mm = 44.01g/mol

= 0.04401 Kg/mol

T= 100.0 C

= (100.0+273) K

= 373 K

vrms = sqrt (3"\times" R"\times" "\frac{T}{M}" )

here : T is temperature in Kelvin

and M is molar mass in Kg/mol

So,

vrms = sqrt(3"\times" 8.314"\times" "\frac{373.0}{0.04401}" )

= 459.8 m/s

Answer: 459.8 m/s